OpenStudy (anonymous):

Find the distance between the points given. (-3, -4) and (0, 0) -5 √(22) 5

OpenStudy (uri):

The formula of distance will be applied : http://geometry.jdeer.com/Images/c1/distance%20formula.gif And, (-3,-4) =(x1 -y1) (0,0)=(x2-y2)

OpenStudy (anonymous):

what do i do to get the answer then?

OpenStudy (uri):

Put the values in that formula

terenzreignz (terenzreignz):

Distance formula? :D \[\Large d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\]

OpenStudy (uri):

Gave the pic already,don't like to type xD

terenzreignz (terenzreignz):

LaTeX is awsomer :3

OpenStudy (uri):

if only i knew how it works.

OpenStudy (anonymous):

d= 0- (-3)^2 + 0- (-4)^2 ?

OpenStudy (anonymous):

d=25 ?

OpenStudy (anonymous):

ohhhh the square root of 25=d so is 5 the correct answer?

OpenStudy (uri):

yes 5 is the answer :)

OpenStudy (anonymous):

Thank you! ^_^

OpenStudy (uri):

You're welcome :D

OpenStudy (anonymous):

#your new fan :D

OpenStudy (uri):

I'm your first fan :D

OpenStudy (anonymous):

Whoop whoop! ^_^